OpenStudy (anonymous):

what meant by with multiplicity 2 ??

OpenStudy (unklerhaukus):

2 has a multiplicity of 2 equal to 1 because\[2=2^1\] 4 has a multiplicity of 2 equal to 2 because\[4=2\times2\] 16 has a multiplicity of 2 equal to 4 because\[16=2\times 2\times2\times2=2^4\]